OpenStudy (loser66):

How to solve it? find a basis of C^2 such that the matrix with respect to that basis is triangular \[\left[\begin{matrix}1&1\\1&0\end{matrix}\right]\] Please, help

OpenStudy (loser66):

@tkhunny

OpenStudy (loser66):

eigenvalues: \(\lambda_1 = \dfrac{1+\sqrt{5}}{2}\) and \(\lambda_2 = \dfrac{1-\sqrt{5}}{2}\)

OpenStudy (loser66):

but they are so ugly eigenvectors so that I can't step up finding out the triangular matrix A with respect to that basis

OpenStudy (kainui):

Haha, well I'm not entirely sure what you're looking for here. Are you just trying to diagonalize this matrix?

OpenStudy (loser66):

if it 's just diagonalize, I am get A for the course. hahahaha... let me attach one which I can get the result for you to know what I am doing. :)

OpenStudy (kainui):

Sure, I am pretty good with change of basis and transformation kind of stuff, I think we can probably figure this out together.

OpenStudy (loser66):

thanks for being here :)

OpenStudy (loser66):

here is another one

OpenStudy (kainui):

I am sort of messed up I guess after finding the first eigenvector for 56.1 I am sort of lost with what's going on I'm afraid.

OpenStudy (loser66):

this is lecture from the book and the exercise is on the last page

OpenStudy (loser66):

Thanks for support. need sleep now.